OpenStudy (osgood):

Help Will Medal!!!! Use part = percent × whole to find out the percent of change between 36 and 25

OpenStudy (anonymous):

This should be in chemistry... @Osgood

OpenStudy (osgood):

I already know to subtract 25 from 36 but what next? what is the answer?

OpenStudy (osgood):

No this is on math in school

OpenStudy (xmissalycatx):

To calculate the percentage increase: First: work out the difference (increase) between the two numbers you are comparing. Then: divide the increase by the original number and multiply the answer by 100. If your answer is a negative number then this is a percentage decrease.

OpenStudy (osgood):

so the answer is 44% if i understand you right @xMissAlyCatx

OpenStudy (xmissalycatx):

36 - 25 and then divide it by 36 (that's your original #). Then multiply that by 100! :) Can you show me the work for what you did so I can understand your process?

OpenStudy (osgood):

oh I divided it by 25

OpenStudy (osgood):

11/36=.36*100=36% ?

OpenStudy (osgood):

@xMissAlyCatx

OpenStudy (xmissalycatx):

If you divide a number by another number then multiply it, all you will get the original number you had before.. what I mean by that is if you have 10 and you divide it by 5, then you multiply it by 5, you'd get the same answer.

OpenStudy (osgood):

? Was I right or not I do not understand what you said

OpenStudy (anonymous):

3.14159265358979323...

OpenStudy (xmissalycatx):

36 - 25 = 11 11 / 36 = .305....... .305...... (100) = 30.5.....

OpenStudy (osgood):

So 30.5%

OpenStudy (osgood):

@xMissAlyCatx As you can tell i am terrible at math

OpenStudy (osgood):

some one tell me if im right

OpenStudy (anonymous):

I would say it is 30.5% or approximately 31%

OpenStudy (osgood):

So I did my math right?

OpenStudy (anonymous):

According to the steps that were said then yes :)

OpenStudy (osgood):

THANK YOU I gave you a medal

OpenStudy (anonymous):

Thanks :)

OpenStudy (xmissalycatx):

But... I showed you how to do it and the answer was on there. xD

OpenStudy (osgood):

Lol